# Topological flow data analysis for transient flow patterns: a graph-based approach

### v1

@@ 3.1.1. Transition of streamline topology
Considering the lid-driven cavity flow at $Re=14000$, we describe how the topological structures of streamline patterns change by utilising COT representations. While the flow pattern evolves continuously in time, their topological structures remain unchanged for a certain period since they are robust under small continuous perturbations. By TFDA, we show that the evolution of the flow patterns is decomposed into a discrete transition among several topological equivalence classes of flow patterns with the same COT representations. In this analysis, we sample the streamline data every $10$ steps after $t=T_{s}$, i.e., at $t=t_{k}\equiv T_{s}+10(k+1)\Delta t$ for $k=0,1,\cdots.$
@@
Figure 7. Transition of the streamline patterns in the lid-driven cavity flow from $t=t_{140}$ to $t=t_{580}$. The Reynolds number is $Re=14000$.
@@
Figure 7 shows snapshots of the evolution of the flow from $t=t_{140}$ to $t=t_{580}$, in which we observe five topologically different flow patterns whose COT representations are provided in Table 2. The COT representation is uniquely provided to flow patterns with the same topological structure, and this state continues for a certain period in the evolution. Hence, each COT representation is associated with a duration as shown in Table 2. For instance, the flow patterns from $t=t_{0}$ to $t=t_{277}$ have the same topological structure as that for Figures 7(a), so the COT representation $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+}))$ remains unchanged in this period. At $t=t_{278}$, the topological structure of the flow pattern transitions to the one with the same COT representation as that for Figure 7(b), i.e., $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+}))$, and it continues until $t=t_{318}$. Similarly, we have the duration associated with the COT representations. Hence, the flow patterns in Figure 7 are regarded as the representative elements of the topologically equivalent classes, and the evolution of the flow patterns is thus reduced to a discrete transition between the COT representations.
Figure time COT duration
7(a) $t=t_{140}$ $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+}))$ $[t_{0},t_{277}]$
7(b) $t=t_{300}$ $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+}),c_{+}^{2}(\sigma_{+}))$ $[t_{278},t_{318}]$
7(c) $t=t_{360}$ $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+}))$ $[t_{319},t_{397}]$
7(d) $t=t_{400}$ $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+}))$ $[t_{398},t_{424}]$
7(e) $t=440$ $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-})),c_{+}^{2}(\sigma_{+}))$ $[t_{425},t_{497}]$
7(f) $t=580$ $\beta_{\emptyset-}(\sigma_{-},c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+}))$ $[t_{498},t_{580}]$
@@
Owing to the uniqueness, using the COT symbols contained in the COT representations, we can describe how the topological structure of the flow changes in these periods. Since all the COT representations start from the same COT symbol $\beta_{\emptyset-}$ expressing that the root structure is a large clockwise rotational flow, we pay attention to the local flow structure represented by the sequence of COT symbols inside the parenthesis of $\beta_{\emptyset-}$. The first COT symbol in the sequence is $\sigma_{-}$ in all cases, indicating the existence of an isolated elliptic fixed point at the centre of the large rotational flow. In the COT representation for Figure 7(a) at $t=t_{140}$, the three COT symbols $c_{+}^{i}(\sigma_{+})$ for $i=0,1,2$ follows $\sigma_{-}$. Each symbol represents an anti-clockwise recirculating flow region enclosed by a boundary saddle connection at the corner of the cavity. In Figure 7(a), these elliptic centres in the recirculating flow domains are plotted using markers $+$, $\times$, $\star$, and $\bullet$. At $t=t_{300}$, another $c_{+}^{0}(\sigma_{+})$ is added to the COT representation for the pattern of Figure 7(b), indicating that a new boundary saddle connection appears at the top left corner. The pair of the COT symbols $c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})$ in the COT representation at $t=t_{300}$ is then replaced by $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$ at $t=t_{360}$. This means that the two boundary saddle connections merge and transition to a local orbit structure of a boundary saddle connection containing a figure-eight orbit structure as observed at the top left corner in Figure 7(c). Afterwards, $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$ is reduced to a simple $c_{+}^{0}(\sigma_{+})$ in the COT representation at $t=t_{440}$. Hence, the figure-eight orbit structure at the top left corner disappears and transforms into a simple boundary saddle connection as in Figure 7(d). The COT symbol $c_{+}^{1}(\sigma_{+})$ at $t=t_{300}$ similarly changes as $c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\rightarrow c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\rightarrow c_{+}^{1}(\sigma_{+})$ from $t=t_{360}$ to $t=t_{580}$. This also indicates that the local orbit structure at the bottom left corner is subject to the same topological transitions as that at the top left corner. After these transitions occur, the COT representation at $t=t_{580}$ finally returns the same one at $t=t_{140}$.
@@
Figure 8. Generation of a local orbit structure represented by $c_{+}(\sigma_{+})$ through a pinching transition.
@@
According to Sakajo and Yokoyama (2015), by observing the changes in the COT symbols, we can theoretically describe generic topological transitions between local flow patterns and specify marginal structurally unstable flow patterns between them. The generation process of a boundary saddle connection $c_{+}^{0}(\sigma_{+})$ at the top left corner from $t=t_{140}$ to $t=t_{300}$ is schematically provided in Figure 8. From Figure 8(a) to Figure 8(c), a degenerate boundary stagnation point appears on the boundary as in Figure 8(b) and it divides into two regular boundary saddles. This transition is called a pinching transition. Note that the flow of Figure 8(b) is hardly observed in the numerical simulation since it is structurally unstable. On the other hand, the transition of the boundary saddle connection at the top left corner from $t=t_{140}$ to $t=t_{360}$ is expressed as the following changes in the COT symbols.
@@
Based on this, we explain the topological transition of the local flow patterns at the top left corner in this period. See Figure 9. From $t=t_{140}$ to $t=t_{300}$, the two boundary saddle connections represented by $c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})$ in Figure 9(a) approach and a structurally unstable flow pattern with heteroclinic connections emerges as in Figure 9(b). Then a slight perturbation instantaneously gives rise to the transition to a structurally stable pattern with a figure-eight orbit structure as in Figure 9(c). This transition is called a heteroclinic transition. The local orbit structure after the heteroclinic transition is represented by the sequence of the COT symbol $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$, indicating that a small clockwise recirculating flow region $c_{-}(\sigma_{-})$ exists inside the boundary saddle connection $c_{+}^{0}$; From $t=t_{300}$ to $t=t_{360}$, we need two more steps to complete the transition. The elliptic centres in the figure-eight pattern merge, and the flow pattern transfers to that of Figure 9(d) which is represented by $c_{+}^{0}(\sigma_{+},c_{-}(\sigma_{-}))$. This structure then transfers to a simple local orbit structure of $c_{+}^{0}(\sigma_{+})$ in Figure 9(f) through a marginal pattern in Figure 9(e) having a degenerate saddle on the boundary. This is another pinching transition. Hence, the theory concludes that the change in the COT symbols of (8).
Although we expect the existence of the structurally stable local flow pattern represented by $c_{+}^{0}(\sigma_{+},c_{-}(\sigma_{-}))$ as in Figure 9(d) in theory, the height difference of the Hamiltonian function for the internal orbit structure $c_{-}(\sigma_{-})$ is so small that psiclone can cut it off in practice due to the threshold parameter $\varepsilon$. There occurs the same transition of the anti-clockwise recirculating region represented by $c_{+}^{1}$ at the bottom left corner from $t=t_{400}$ to $t=t_{580}$.

### v2

@@ 3.1. Transition of streamline topology
Considering the lid-driven cavity flow at $Re=14000$, we describe how the topological structures of the streamline patterns change using COT representations. Although the flow pattern evolves continuously over time, their topological structures remain unchanged for a certain period, since they are robust under small continuous perturbations. By TFDA, we show that the evolution of the flow patterns is decomposed into a discrete transition among several topological equivalence classes of flow patterns with the same COT representations. In this analysis, we sample the streamfunction data every $10$ step after $t=T_{s}$, that is, at $t=t_{k}\equiv T_{s}+10\times k\Delta t$ for $k=0,1,\ldots,580$.
@@
Figure 8. Transition of the streamline patterns in the lid-driven cavity flow from $t=t_{140}$ to $t=t_{440}$. The Reynolds number is $Re=14000$.
@@
Figure 8 shows snapshots of the evolution of the flow from $t=t_{140}$ to $t=t_{440}$. We observe six topologically different flow patterns whose COT representations are provided in Table 2. The COT representation is unique for flow patterns with the same topological structure, and this state continues for a certain period of evolution. Hence, each COT representation is associated with a duration, as shown in Table 2. For instance, the flow patterns from $t=t_{0}$ to $t=t_{277}$ have the same topological structure as that of Figures 8(a), so the COT representation $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ remains unchanged in this period. At $t=t_{278}$, the topological structure of the flow pattern transitions to the one with the same COT representation as that of Figure 8(b), i.e. $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$. The same COT representation continues until $t=t_{316}$. Hence, the flow patterns in Figure 8 are considered representative elements of each topologically equivalent class, and therefore the evolution of the flow patterns is reduced to a discrete transition between the COT representations. We note that a tiny boundary saddle separatrix always exists in the top right corner of each panel in Figure 8. However, this structure is disregarded by psiclone. Since the streamfunction values of the boundary separatrix and of the elliptic centre inside are close enough to be less than the threshold $\varepsilon=5.0\times 10^{-4}$, the structure is considered a topologically inessential structure, as explained in the previous section. In addition, we observe small boundary saddle separatrices on the left wall in $t=t_{140}$ (Figure 8(a)) and on the bottom wall in $t=t_{300}$ (Figure 8(b)). They are also considered small structures at these times. However, at later times, the difference in the level values of the streamfunction for them becomes larger than the threshold $\varepsilon$ and they are eventually recognised as boundary saddle separatrices as shown in Figures 8(b,c).
ID Figure time COT duration 1 8(a) $t_{140}$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{0},t_{277}]$ 2 8(b) $t_{300}$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{278},t_{316}]$ 3 8(c) $t_{317}$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{317}$ 4 8(d) $t_{360}$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{318},t_{397}]$ 5 8(e) $t_{400}$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{398},t_{424}]$ 6 8(f) $t_{440}$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{425},t_{497}] - - -$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ $t_{498},t_{580}]$
@@
Based on the COT representations in Table 2, we describe how the topological structure of the flow changes during this period. As mentioned in the previous section, psiclone can compute the COT representation for a given streamfunction without numerically determining centres, saddles, or separatrices. On the other hand, in Figure 8, we plot critical points using markers when the difference in the level values of the streamfunction for them becomes greater than the threshold $\varepsilon$, and we plot separatrices for reference. All COT representations start from the same COT symbol $\beta_{\emptyset-}$ whose first component is $\sigma_{-}$ indicating the existence of an isolated elliptic fixed point at the centre of the large clockwise rotational flow in the cavity. Hence, we pay attention to the local flow structure represented by the sequence of COT symbols inside the curly brackets of the COT symbol for $\beta_{\emptyset-}$. In the COT representation for Figure 8(a) at $t=t_{140}$, there are the three COT symbols $c_{+}^{i}(\sigma_{+})$, $i=0,1,2$ inside the curly brackets. Each symbol represents an anticlockwise recirculating flow region enclosed by a boundary saddle connection in the corner of the cavity. Here, we focus on the change in the COT symbols for $c_{+}^{0}$. At $t=t_{300}$, another $c_{+}^{0}(\sigma_{+})$ is added to the COT representation for the pattern of Figure 8(b), indicating that a new boundary saddle connection appears on the left wall in the region of the top left corner. With a brief fluctuation in $t=t_{317}$ with respect to the bottom left corner as shown in Figure 8(c), the pair of COT symbols $c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})$ in $t=t_{300}$ is then replaced by $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$ in $t=t_{360}$. This means that the two boundary saddle connections merge and transition to a local orbit structure of a boundary saddle connection that contains a figure-eight orbit structure as observed in the top left corner of Figure 8(d). Subsequently, $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$ is reduced to a simple $c_{+}^{0}(\sigma_{+})$ at $t=t_{400}$. Hence, the figure-eight orbit structure disappears and transforms into a simple boundary saddle connection as in Figure 8(d). The COT symbol $c_{+}^{1}(\sigma_{+})$ for the bottom left corner eddy at $t=t_{300}$ changes similarly as $c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\rightarrow c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\rightarrow c_{+}^{1}(\sigma_{+})$ from $t=t_{317}$ to $t=t_{580}$. This also indicates that the local orbit structure in the bottom left corner is subject to the same topological transitions as that in the top left corner. After these transitions occur, the COT representation finally returns the same COT representation at $t=t_{498}$ as that of $t=t_{140}$, and it continues until $t=t_{580}$.
@@
Figure 9. Generation of a local orbit structure represented by $c_{+}(\sigma_{+})$ on the left wall through a pinching transition.
@@
Sakajo and Yokoyama (2015) proposed a theory to interpolate topological pattern transitions that can occur from changes in COT symbols. Based on this theory, it is possible to describe an achievable pathway in the region of the top left corner even if the time series data do not contain intermediate patterns of the transition. First, the generation of a boundary saddle connection $c_{+}^{0}(\sigma_{+})$ on the left wall in the region of the top left corner from $t=t_{140}$ to $t=t_{300}$ is schematically provided in Figure 9. From Figure 9(a) to Figure 9(c), a degenerate boundary stagnation point appears on the left wall as in Figure 9(b) and divides into two regular boundary saddles. This transition is called a pinching transition. The flow of Figure 9(b) is hardly observed in the numerical simulation, since it is structurally unstable. Next, the transition of the boundary saddle connection in the top left corner from $t=t_{300}$ to $t=t_{400}$ is expressed as the following changes in the COT symbols.
@@
An interpolated transition pathway is shown schematically in Figure 10. From $t=t_{300}$ to $t=t_{360}$, the two boundary saddle connections represented by $c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})$ in Figure 10(a) approach, and a self-connected saddle connection between them (Figure 10(c)) emerges through a pinching transition (Figure 10(b)). These local orbit structures merge to form a structurally unstable pattern with heteroclinic connections (Figure 10(d)). Note that in the present TFDA, no COT representation corresponding to the pattern of Figure 10(c) is observed. This indicates that this transition occurs very quickly, and so streamfunction with this pattern was not contained in the time-series data. However, even in this situation, the transition theory can propose this route as one of the possible transitions. A slight perturbation of Figure 10(d) instantaneously gives rise to the transition to the same structurally stable pattern with a figure-eight orbit structure as the pattern at $t=t_{360}$ in Figure 10(e). This transition is called a heteroclinic transition. The local orbit structure after the heteroclinic transition is represented by COT symbols $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$, indicating that a small clockwise recirculating flow region $c_{-}(\sigma_{-})$ on the left wall is enclosed by the anticlockwise saddle connection $c_{+}^{0}$. From $t=t_{360}$ to $t=t_{400}$, we need two more steps to complete the transition. The small clockwise recirculating flow $c_{-}(\sigma_{-})$ disappears due to the reverse pinching transition on the left wall (Figure 10(f)). Consequently, this produces a structurally stable flow pattern as in Figure 10(g). This flow pattern is also not detected by TFDA because it persists for a short time in the evolution. Finally, the elliptic centre and the saddle in the figure-eight pattern merge by a pinching transition (Figure 10(h)), and the flow pattern transfers to a simple local orbit structure of $c_{+}^{0}(\sigma_{+})$ like Figure 10(i) at $t=t_{400}$.